### v1

@@ 5.3.1 Comparison of Gen 1A, 1B, and 1C
Table 6: Comparison of GEN 1C’s Performance under Different Initial Learning Rates
Metric Lr = $1\times 10^{-4}$ Lr = $7.5\times 10^{-5}$
Epochs val_loss TSR(%) Epochs val_loss TSR(%)
Early 65 0.799 73.958
Mid 150 0.752 72.917 155 0.764 77.083
Late 190 0.746 72.917 200 0.757 75.000
[…]
5.3.2 Comparison of Gen 2A and 2B
Based on the experience gained from Gen 1, an initial learning rate of $7.5\times 10^{-5}$ will be chosen for Gen 2, with a maximum of 155 epochs of training, to achieve a balance between model performance and training duration. The result is demonstrated in Table 7.
In Gen 2A, increasing the number of additional scenes without increasing the dataset size did not have a positive effect on model performance, as the TSR of 63.542% was 13.54 percentage points lower than the best result achieved by Gen 1C. The scene composition of Gen 2B has 16 more scenes compared to Gen 1C, ensuring that the additional scenes have sufficient parking route data. The model based on Gen 2B achieves the same TSR (77.083%) as that of Gen 1C. Due to the shorter straight-line distance before braking in Gen 2, the model performs better on parking slots near the edge of the parking lot. In Gen 1C, the TSR for slots 2-1 and 3-1 is only 16.667%, whereas in Gen 2B, the TSR for these slots is 66.667% and 50%, respectively.
Table 7: Comparison of Gen 2A and Gen 2B’s Performance
Metric Gen 2A Gen 2B
Epochs 150 130
val_loss 0.9550 0.8089
TSR(%) 63.542 77.083
[…]
00footnotetext: The initial learning rate is set as $7.5\times 10^{-5}$ during training on both datasets.
5.3.3 Details of Gen 2B testing
The preliminary test result of model on Gen 2B is illustrated in Table 8. The experiments consist of 96 tests in 16 different scenes. By observing the real-time video of the model’s inference and control of the vehicle, it is found that the vehicle’s speed is too high, leading to unintended movement towards non-target parking slots or even collisions. To address this issue, the speed for both gear 1 and reverse (R) was limited to 7.5 km/h. Additionally, a closed-loop experiment consisting of 384 test cases was conducted to evaluate the effectiveness of this adjustment. The result in Table 9 demonstrates that the model trained based on Gen 2B demonstrates robust performance, achieving a TSR of 85.156%. Compared to the original speed limit set by the author, the stricter speed limit resulted in an increase of 1.417 seconds in APT. However, this adjustment led to a 6.69% reduction in APE and a 14.76% decrease in AOE. More importantly, CR decreased from 9.375% to 2.865%, improving safety.
@@ 5.3.3 Details of Gen 2B testing
Table 8: Closed-loop results including 96 test cases for model trained on Gen 2B, with the speed limit set as default.
TaskIdx TSR TFR NTSR NTFR CR OR TR APE AOE APT AIT
@@
[…]
Table 9: Closed-loop results including 384 test cases for model trained on Gen 2B, with the speed limit set as 7.5 km/h.
TaskIdx TSR TFR NTSR NTFR CR OR TR APE AOE APT AIT
@@
[…]
5.3.4 Parking performance compared with baseline and original expert
According to Table 10, the model trained on the new dataset Gen 2B can essentially reproduce the performance of the baseline. Although TSR is 6.25 percentage points lower and APT increases by 6.36 seconds, APE decreases by 20%, and AOE decreases by 60.9%. Meanwhile, the new expert shows a significant improvement in accuracy compared to the original expert. Although APT increases by 1.99 seconds, APE decreases by 43.5% and AOE decreases by 90.0%.
Table 10: Performance comparison with baseline and original expert
@@ 5.3.4 Parking performance compared with baseline and original expert
[…]
6 Conclusion
In this work we created an open-source dataset for E2E Parking and achieved a Target Success Rate comparable with the original authors. Moreover, the model trained on the dataset shows an enhanced parking precision with lower Average Position Error and Average Orientation Error. The iterative process of dataset construction provides insights for relevant researchers on how to optimize the dataset and improve training and testing results through methods such as parking operation strategies, scene selection, learning rate adjustment, and speed limitations.

### v2

@@ V-C3 Details of Gen 1C testing
Table VIII demonstrates the results of the best model trained on Gen 1C. With the help of the newly added parking trajectories featuring shadow projected on the target slot, the new model shows a significant improvement on slot 3-9, with the TSR increasing from 0 to 66.667%. This improvement in slot 3-9 contributes to a 5.208 percentage point increase in the overall average TSR compared to Gen 1B. TSRs for other slots show no notable changes, and slots 2-1 and 3-1 still maintain very low TSR levels, which is the primary focus during creation of dataset Gen 2.
@@
TABLE VIII: Closed-loop results including 96 test cases for the best model trained on Gen 1C
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
2-1 16.667 0.000 33.333 0.000 16.667 33.333 0.000 0.383 0.363 26.800 0.077
2-3 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.304 0.258 22.887 0.077
2-5 50.000 0.000 0.000 0.000 0.000 0.000 50.000 0.377 0.608 23.989 0.077
2-7 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.352 0.284 23.061 0.077
2-9 83.333 0.000 16.667 0.000 0.000 0.000 0.000 0.462 0.338 22.300 0.077
2-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.410 0.464 22.811 0.078
2-13 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.354 0.988 22.122 0.077
2-15 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.345 1.093 23.206 0.077
3-1 16.667 0.000 0.000 16.667 16.667 33.333 16.667 0.290 1.019 23.333 0.077
3-3 66.667 0.000 16.667 0.000 16.667 0.000 0.000 0.240 0.451 24.283 0.077
3-5 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.328 0.211 23.550 0.078
3-7 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.344 0.707 21.356 0.078
3-9 66.667 0.000 16.667 0.000 16.667 0.000 0.000 0.264 0.984 23.242 0.077
3-11 66.667 0.000 33.333 0.000 0.000 0.000 0.000 0.240 0.361 23.775 0.077
3-13 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.187 0.307 25.693 0.077
3-15 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.176 0.339 24.506 0.077
Avg 77.083 0.000 7.292 1.042 4.167 4.167 6.250 0.316 0.548 23.557 0.077
V-C4 Details of Gen 2A testing
Based on the experience gained from Gen 1, an initial learning rate of $7.5\times 10^{-5}$ will be chosen for Gen 2, with a maximum of 155 epochs of training, to achieve a balance between model performance and training duration. The result is demonstrated in Table XIV.
In Gen 2A, increasing the number of additional scenes without increasing the dataset size did not have a positive effect on model performance, as the TSR of 63.542% was 13.54 percentage points lower than the best result achieved by Gen 1C.
Table IX demonstrates the test results of the model trained on Gen 2A. Compared to Gen 1C, the new model performs better with slot 2-1, 2-5 and 3-1, which indicates that the shortened forward-driving distance before reversing is helpful in handling the scenarios near the edge of the parking lot. However, there are 10 slots where the TSR drops with the new model, which proves that t is not adequate to generate only one trajectory for a single parking slot in a given scenario and it is unwise to increase the number of parking slot scenarios while keeping the total number of trajectories unchanged. Overly complicated scenario setups are not an effective solution for optimizing the dataset.
TABLE IX: Closed-loop results including 96 test cases for the model trained on Gen 2A
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
2-1 50.000 0.000 0.000 0.000 16.667 33.333 0.000 0.247 0.594 19.111 0.077
2-3 50.000 16.667 16.667 16.667 0.000 0.000 0.000 0.219 0.193 20.056 0.078
2-5 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.307 0.775 20.700 0.080
2-7 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.264 0.524 20.411 0.079
2-9 33.333 0.000 50.000 16.667 0.000 0.000 0.000 0.113 0.214 20.333 0.079
2-11 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.183 0.785 21.533 0.080
2-13 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.222 0.709 20.380 0.079
2-15 50.000 0.000 0.000 0.000 33.333 0.000 16.667 0.140 0.534 20.767 0.079
3-1 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.319 0.708 21.087 0.078
3-3 66.667 0.000 16.667 0.000 16.667 0.000 0.000 0.175 0.217 20.758 0.079
3-5 66.667 0.000 0.000 0.000 33.333 0.000 0.000 0.166 0.143 20.692 0.080
3-7 66.667 0.000 16.667 0.000 16.667 0.000 0.000 0.233 0.546 21.492 0.081
3-9 16.667 0.000 16.667 0.000 33.333 0.000 33.333 0.121 0.031 22.833 0.079
3-11 33.333 0.000 50.000 0.000 0.000 0.000 16.667 0.320 1.603 21.483 0.078
3-13 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.137 0.245 22.420 0.078
3-15 66.667 0.000 0.000 0.000 0.000 16.667 16.667 0.120 0.148 23.717 0.078
Avg 63.542 1.042 10.417 2.083 13.542 3.125 6.250 0.205 0.498 21.111 0.079
V-C5 Details of Gen 2B testing
The scene composition of Gen 2B has 16 more scenes compared to Gen 1C, ensuring that each of the additional scenes has sufficient parking route data. The model based on Gen 2B achieves the same TSR (77.083%) as that of Gen 1C. Due to the shorter straight-line distance before braking in Gen 2, the model performs better on parking slots near the edge of the parking lot. In Gen 1C, the TSR for slots 2-1 and 3-1 is only 16.667%, whereas in Gen 2B, the TSR for these slots is 66.667% and 50%, respectively.
The preliminary test result of model on Gen 2B is illustrated in Table X. The experiments consist of 96 tests in 16 different scenes. By observing the real-time video of the model’s inference and control of the vehicle, it is found that the vehicle’s speed is too high, leading to unintended movement towards non-target parking slots or even collisions.
To address this issue, the speed for both gear 1 and reverse (R) was limited to 7.5 km/h. Additionally, a closed-loop experiment consisting of 384 test cases was conducted to evaluate the effectiveness of this adjustment. The result in Table XI demonstrates that the model trained based on Gen 2B has a robust performance, achieving a TSR of 85.156%. The model performs better in 7 tasks and maintains the same TSR in 3 tasks. Compared to the original speed limit set by the author, the stricter speed limit resulted in an increase of 1.417 seconds in APT. However, this adjustment led to a 6.69% reduction in APE and a 14.76% decrease in AOE. More importantly, CR decreased from 9.375% to 2.865%, improving safety.
TABLE X: Closed-loop results including 96 test cases for the model trained on Gen 2B, with the speed limit set as default.
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
@@ V-C5 Details of Gen 2B testing
TABLE XI: Closed-loop results including 384 test cases for model trained on Gen 2B, with the speed limit set as 7.5 km/h.
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
@@
V-C6 Parking performance compared with baseline and original expert
According to Table XII, the model trained on the new dataset Gen 2B can essentially reproduce the performance of the baseline. Although TSR is 6.25 percentage points lower and APT increases by 6.36 seconds, APE decreases by 20%, and AOE decreases by 60.9%. Meanwhile, the new expert shows a significant improvement in accuracy compared to the original expert. Although APT increases by 1.99 seconds, APE decreases by 43.5% and AOE decreases by 90.0%.
TABLE XII: Performance comparison with baseline and original expert
@@ V-C6 Parking performance compared with baseline and original expert
TABLE XIII: Comparison of GEN 1C’s Performance under Different Initial Learning Rates
Metric Lr = $1\times 10^{-4}$ Lr = $7.5\times 10^{-5}$
Epochs val_loss TSR(%) Epochs val_loss TSR(%)
Early 65 0.799 73.958
Mid 150 0.752 72.917 155 0.764 77.083
Late 190 0.746 72.917 200 0.757 75.000
TABLE XIV: Comparison of Gen 2A and Gen 2B’s Performance
Metric Gen 2A Gen 2B
Epochs 150 130
val_loss 0.9550 0.8089
TSR(%) 63.542 77.083
44footnotetext: The initial learning rate is set as $7.5\times 10^{-5}$ during training on both datasets.
VI Conclusion
In this work we created an open-source dataset for E2E Parking and achieved a Target Success Rate outperforming the prior work. Moreover, the model trained on the dataset shows an enhanced parking precision with lower Average Position Error and Average Orientation Error. The iterative process of dataset construction provides insights for relevant researchers on how to optimize the dataset and improve training and testing results through methods such as parking operation strategies, scene selection, learning rate adjustment, and speed limitations.